OpenStudy (anonymous):

Given that Sin A = 4/5, 0 < A < π/2 and Cos B = -12/13, π< B < 3π/2, find Sin (A+B) choices: a.) -56/65 b.) 56/65 c.) 63/16 d.) -63/16

OpenStudy (raden):

use this formula Sin (A+B) = sin(A)cos(B) + cos(A)sin(B)

OpenStudy (anonymous):

thank so much :)

OpenStudy (raden):

now find cos(A) if given sin(A) = 4/5 use the identity : cos^2 (A) = 1 - sin^2 (A) cos^2 (A) = 1 - (4/5)^2 cos^2 (A) = 1 - 16/25 cos^2 (A) = 9/25 cos(A) = 3/5 (notice that A in the 1st quadrant, so cos be positive)

OpenStudy (raden):

then find sin(B). given Cos B = -12/13, sin^2 (B) = 1 - cos^2 (B) sin^2 (B) = 1- (-12/13)^2 sin^2 (B) = 1- 144/169 sin^2 (B) = 25/169 sin(B) = -5/13 (see that B in the 3rd quadrant, so sin be negative)

OpenStudy (raden):

finally, plugin all values to formula sin(a+b) sin (A+B) = sin(A)cos(B) + cos(A)sin(B) sin (A+B) = 4/5 * (-12/13) + 3/5 * (-5/13) = ... simplify it to get your answer

OpenStudy (anonymous):

answer is -113/65 ?

OpenStudy (raden):

how do you get -113 ?

OpenStudy (anonymous):

-63 / 65 :) is the correct answer :) thank you my friend

OpenStudy (raden):

case is closed :)

OpenStudy (raden):

welcome

OpenStudy (anonymous):

:D thank you ..help me again i have so many problems here ahahaha. Given that Sin A = 4/5, 0 < A < π/2 and Cos B = -12/13, π< B < 3π/2, find Tan (A+B) .. this time its tan(A+B) :)

OpenStudy (raden):

well, it like before we have sin(A) = 4/5, cos(A) = 3/5, sin(B) = -5/13, and cos(B) = -12/13 now find tan(A) and tan(B) respectively is ... see tan = sin/cos, so tan(A) = sin(A)/cos(A) = 4/5)/(3/5) = 4/3 (tan in the 1st quadrant is +) tan(b) = sin(B)/cos(B) = (5/13)/(12/13) = 5/12 (tan in the 3rd quadrant is +)

OpenStudy (raden):

the question is tan(A+B). use the identity : tan(A+B) = (tan(A) + tan(B))/(1 - tan(A)tan(B)) so, tan(A+B) = (4/3 + 5/12)/(1 - 4/3 * 5/12) tan(A+B) = .... ?

OpenStudy (anonymous):

63/16 is the answer :) thank you again

OpenStudy (anonymous):

Given that Sin A = 4/5, 0 < A < π/2 and Cos B = -12/13, π< B < 3π/2, find Cos 2A